OpenStudy (anonymous):

what is the horizontal change between the points (-2, 5) and (4, 10)?

OpenStudy (anonymous):

Just subtract the x-values. 4 - (-2) = 4 + 2 = ?

OpenStudy (anonymous):

4- -2

OpenStudy (mertsj):

Well think of a number line. How far is it from -2 to 4?

OpenStudy (unklerhaukus):